# MARS: Reinforcing Multi-Agent Reasoning of LLMs through Self-Play in Strategic Games

### v1

@@ 6 Discussion
In this work, we introduced MARS, a framework that uses self-play in a diverse suite of cooperative and competitive games to cultivate multi-agent capabilities in LLMs. Our findings demonstrate that skills honed in this curriculum directly transfer to and enhance performance in multi-agent systems, establishing self-play as a scalable, autonomous paradigm for training LLM agents.
Despite its promising results, MARS has several limitations that present avenues for future work. First, our current study is confined to two-player games with relatively constrained action spaces. Scaling this approach to more complex, N-player environments will pose significant challenges in credit assignment and strategic modeling. Moreover, while we demonstrate the transferability of learning in strategic games, a particularly compelling direction is the development of more complex "social sandboxes" beyond classic games, designed to train agents in more practical domains like software engineering or scientific discovery.
Ultimately, our work provides strong evidence that the principles of self-play are a powerful engine for progress, paving the way for the next generation of sophisticated, socially intelligent agents.
@@
All code, model checkpoints, and training scripts required to reproduce the findings of this paper are publicly available at https://github.com/thu-nics/MARS. The repository includes all necessary configurations to replicate our key experiments.
@@
This work adheres to the ICLR Code of Ethics. The primary ethical consideration of our work is the dual-use nature of enhancing LLM agent capabilities. While our goal is to foster beneficial multi-agent skills like collaboration and competition, the resulting models could potentially be applied to malicious ends. We release our code and models to the research community to encourage further study into the safety and alignment of more capable agents.
@@ Appendix B Implementation details
Our implementation of MARS is built upon ROLL (Wang et al., 2025), a robust open-source codebase for reinforcement learning from LLM feedback. ROLL’s native support for agentic, multi-turn rollouts provided a strong foundation for our system. To achieve high performance, ROLL leverages vLLM for efficient inference during the rollout phase (Kwon et al., 2023) and is built on Megatron-LM for distributed training (Shoeybi et al., 2019). All game environments were implemented using OpenSpiel (Lanctot et al., 2019), ensuring the correctness and standardization of game logic.
@@
We use Qwen-4B as our base model for all experiments. The training is conducted in a fully online manner, where self-play trajectories are immediately used for policy updates. Specialist agents are trained with a batch size of 128 trajectories, while the generalist agent uses a batch size of 384 (128 from each game). All models were trained for a total of 200 optimization steps. For optimization, we employ the Adam optimizer (Kingma, 2014) with a cosine annealing learning rate schedule, which warms up for 10 steps to a peak of $1\times 10^{-6}$ before decaying.
@@
All training was conducted on a single machine equipped with 8 NVIDIA H100 GPUs.
@@ Appendix C Evaluation details
To rigorously assess game-playing proficiency, we evaluate our agents against a suite of strong, fixed opponents. For adversarial games, this includes:
  • Tic-Tac-Toe and Connect Four: MCTS agents with varying simulation counts (100/1000 for Tic-Tac-Toe, 10/100 for Connect Four) to test against different strengths.
@@
In all adversarial games, models are evaluated as both the first-move and second-move player, measured by the average game return. For the cooperative Hanabi variants, performance is measured by the standard self-play game return. In all game settings, agents are evaluated for 1000 games.
@@
For downstream task evaluation, we test all models on a suite of mathematics and general QA benchmarks. In the standard single-agent setting, we use the evaluation scripts from Qwen2.5-Math (Yang et al., 2024a; b) for MATH500, GSM8K, AQUA-RAT, AIME24, AMC23, and MMLU-STEM, and lm-evaluation-harness (Gao et al., 2024) for GPQA-Diamond. For the multi-agent evaluation, we utilize the MASLab framework (Ye et al., 2025), which integrates multi-agent LLM systems like MAD and AutoGen, for evaluation of all benchmarks.
Across all benchmark evaluations, we maintain consistent generation parameters for a fair comparison: a temperature of 0.6, top-p of 0.95, and top-k of 20. To balance evaluation efficiency with accuracy, we set the maximum output length to 8,192 tokens.
Appendix D Hyperparameters
@@
Appendix E Game observation and prompt
@@ 2. In each of your turns:
Mini Hanabi
For Mini Hanabi, a cooperative imperfect information game, we provide the agent with a unique observation. An agent observes the cards held by all other players but remains unaware of its own hand. The observation also includes the number of remaining information and life tokens, the cards successfully played on the board (fireworks), and the contents of the discard pile. The prompt is highly structured, detailing the game rules, player-specific information, and strict response formatting. For training, we use MiniHanabi, a simplified 2-player version with 2 colors, 2 ranks, 3-card hands, and 3 information/life tokens. For out-of-distribution evaluation, we use a more complex variant, SimpleHanabi, which features 3 colors, 2 ranks, 5-card hands, 8 information tokens and 3 life tokens. For example, an initial prompt details the starting setup of the chosen variant, including all visible cards and token counts, before asking the first player to make the first move (Listing 3).
@@
Appendix F Full results
Table 6: Full results on game-play return in training games
Model TicTacToe KuhnPoker MiniHanabi
MCTS(100 sims) MCTS(1000 sims)
@@ Appendix F Full results
MARS (Tic.) 0.506/-0.358 -0.278/-0.379 -0.119/-0.142 2.019
MARS (Kuhn.) 0.397/-0.490 -0.388/-0.538 -0.107/-0.103 1.799
MARS (Hanabi) 0.438/-0.644 -0.350/-0.702 -0.123/-0.130 2.222
MARS (Multi.) 0.544/-0.419 -0.302/-0.450 -0.114/-0.141 2.173
+ fix opponent
MARS (Tic.) 0.760/-0.161 -0.378/-0.162 -0.146/-0.090 1.397
MARS (Kuhn.) -1.000/-1.000 -1.000/-1.000 -0.114/-0.117 0.000
MARS (Hanabi) - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) 0.492/-0.517 -0.292/-0.567 -0.104/-0.091 1.392
MARS (Kuhn.) 0.267/-0.607 -0.426/-0.625 -0.074/-0.106 1.666
MARS (Hanabi) 0.351/-0.788 -0.400/-0.811 -0.133/-0.084 2.128
+ w/o agent-specific advantage normalization
MARS (Tic.) 0.654/-0.376 -0.213/-0.366 -0.127/-0.126 1.764
MARS (Kuhn.) 0.391/-0.509 -0.446/-0.567 -0.116/-0.109 1.607
MARS (Hanabi) 0.363/-0.732 -0.364/-0.764 -0.119/-0.128 2.100
@@
Model ConnectFour LeducHoldem SimpleHanabi
MCTS(10 sims) MCTS(100 sims)
@@
MARS (Tic.) -0.387/-0.703 -0.858/-0.974 -0.518/-0.702 1.785
MARS (Kuhn.) -0.448/-0.746 -0.874/-0.995 -0.460/-0.327 1.761
MARS (Hanabi) -0.465/-0.885 -0.880/-0.998 -0.736/-0.403 2.036
MARS (Multi.) -0.421/-0.780 -0.871/-0.979 -0.483/-0.487 2.205
+ fix opponent
MARS (Tic.) -0.593/-0.887 -0.939/-0.990 -0.632/-0.616 0.733
MARS (Kuhn.) -1.000/-1.000 -1.000/-1.000 -1.124/-1.003 0.000
MARS (Hanabi) - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) -0.465/-0.754 -0.843/-0.989 -0.622/-0.553 1.143
MARS (Kuhn.) -0.408/-0.783 -0.899/-0.998 -0.789/-0.663 1.379
MARS (Hanabi) -0.495/-0.939 -0.886/-0.998 -0.555/-0.483 1.841
+ w/o agent-specific advantage normalization
MARS (Tic.) -0.492/-0.790 -0.882/-0.983 -0.594/-0.460 1.303
MARS (Kuhn.) -0.460/-0.790 -0.908/-0.992 -0.753/-0.769 1.345
MARS (Hanabi) -0.358/-0.898 -0.902/-0.998 -0.664/-0.389 1.925
@@
MARS (Tic.) 89.10 95.20 46.50 40.00 77.50 57.60 38.89
MARS (Kuhn.) 87.80 94.50 48.40 33.30 72.50 59.30 33.84
MARS (Hanabi) 88.10 94.70 48.00 43.30 65.00 58.90 36.36
MARS (Multi.) 89.90 94.60 52.00 33.30 75.00 59.90 34.85
+ fix opponent
MARS (Tic.) 89.10 94.80 53.90 50.00 72.50 63.00 37.88
MARS (Kuhn.) 86.20 94.00 45.70 33.30 72.50 53.80 33.84
MARS (Hanabi) - - - - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) 88.80 94.80 48.80 40.00 72.50 57.70 36.36
MARS (Kuhn.) 88.60 94.00 48.40 40.00 72.50 57.90 34.34
MARS (Hanabi) 87.80 94.50 48.40 36.70 70.00 59.30 36.87
+ w/o agent-specific advantage normalization
MARS (Tic.) 88.20 94.50 53.10 36.70 77.50 58.20 39.39
MARS (Kuhn.) 88.10 94.80 53.90 40.00 75.00 58.70 35.86
MARS (Hanabi) 88.60 93.90 46.10 40.00 75.00 57.00 38.89
@@
MARS (Tic.) 92.20 96.06 83.07 43.33 82.50 86.76 41.12
MARS (Kuhn.) 91.60 96.21 82.68 40.00 82.50 87.39 41.41
MARS (Hanabi) 91.40 95.60 82.68 43.33 77.50 87.04 38.38
MARS (Multi.) 92.80 95.60 83.86 46.67 80.00 87.36 45.45
+ fix opponent
MARS (Tic.) 94.80 94.54 85.04 53.33 90.00 88.57 42.93
MARS (Kuhn.) 95.00 94.24 86.22 53.33 85.00 88.54 45.45
MARS (Hanabi) - - - - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) 93.80 94.77 85.04 60.00 85.00 88.98 47.98
MARS (Kuhn.) 95.00 94.54 86.22 56.67 90.00 89.05 50.00
MARS (Hanabi) 94.20 94.62 85.43 53.33 90.00 88.51 46.97
+ w/o agent-specific advantage normalization
MARS (Tic.) 94.40 94.62 84.65 56.67 87.50 89.33 46.97
MARS (Kuhn.) 95.00 94.69 84.25 56.67 85.00 89.62 47.47
MARS (Hanabi) 94.20 94.47 87.01 63.33 90.00 88.92 43.94

### v2

@@ 6 Discussion
In this work, we introduced MARSHAL, a framework utilizing self-play in a diverse range cooperative and competitive strategic games to cultivate multi-agent capabilities in LLMs. Our findings demonstrate that skills honed in strategic games directly transfer to enhanced performance in general multi-agent systems, establishing self-play as a scalable paradigm for training LLM agents.
Despite these promising results, limitations remain. First, our study utilizes two-player games as efficient "prototypes" to cultivate foundational skills. While we demonstrate that these constrained settings are sufficient for strong transfer to general tasks, scaling to larger $N$-player environments introduces significantly greater challenges regarding non-stationarity, population diversity, and credit assignment that warrant dedicated future investigation. Additionally, moving beyond classic games to complex "social sandboxes" (e.g., simulated software engineering, collaborative researching) represents a compelling direction for training agents in more practical domains.
Ultimately, our work provides strong evidence that the principles of self-play are a powerful engine for progress, paving the way for the next generation of sophisticated LLM agents.
@@
All code, model checkpoints, and training scripts required to reproduce the findings of this paper are publicly available at https://github.com/thu-nics/MARSHAL. The repository includes all necessary configurations to replicate our key experiments.
@@
This work adheres to the ICLR Code of Ethics. The primary ethical consideration of our work is the dual-use nature of enhancing LLM agent capabilities. While our goal is to foster beneficial multi-agent skills like cooperation and competition, the resulting models could potentially be applied to malicious ends. We release our code and models to the research community to encourage further study into the safety and alignment of more capable agents.
@@ Appendix B Implementation details
Our implementation of MARSHAL is built upon ROLL (wang2025reinforcement), a robust open-source codebase for post-training LLMs with reinforcement learning. ROLL’s native support for agentic, multi-turn rollouts provided a strong foundation for our system. To achieve high performance, ROLL leverages vLLM for efficient inference during the rollout phase (kwon2023efficient) and is built on Megatron-LM for distributed training (megatron-lm). All game environments used in this work were implemented using OpenSpiel (LanctotEtAl2019OpenSpiel) and VS-Bench (xu2025vs), ensuring the correctness and standardization of game logic.
@@
We use Qwen-4B as our base model for all experiments reported in our main results. The training is conducted in a fully online manner, where self-play trajectories are immediately used for policy updates. Specialist agents are trained with a batch size of 128 trajectories, while the generalist agent uses a batch size of 384 (128 from each game). All models were trained for a total of 200 optimization steps. For optimization, we employ the Adam optimizer (kingma2014adam) with a cosine annealing learning rate schedule, which warms up for 10 steps to a peak of $1\times 10^{-6}$ before gradually decaying to 0.
@@
Models were trained on a single server with 8 NVIDIA H100 GPUs.
@@ Appendix C Evaluation details
To rigorously assess game-playing proficiency, we evaluate our agents against a suite of strong, fixed opponents. For competitive games, this includes:
  • Tic-Tac-Toe and Connect Four: two MCTS agents with varying simulation counts (100/1000 for Tic-Tac-Toe, 10/100 for Connect Four) to test against different strengths.
@@
In all competitive games, models are evaluated as both the first-move and second-move player, measured by the average game return. For the cooperative Hanabi variants, performance is measured by the standard self-play game return. In all game settings, agents are evaluated for 1000 games.
For the results on strategic ability reported in the main text, we normalize the scores with respect to the worst-observed performance and the theoretical optimal performance.
@@
For downstream evaluations, we assess all models on a comprehensive suite of mathematics and general QA benchmarks. In the single-agent setting, we utilize evaluation scripts from Qwen2.5-Math (yang2024qwen2; yang2024qwen2_5_math) for MATH500, GSM8K, AQUA-RAT, AIME24, AMC23, and MMLU-STEM, while employing lm-evaluation-harness (eval-harness) for GPQA-Diamond. For multi-agent evaluations, we adopt the MASLab framework (ye2025maslab) across all benchmarks. To balance evaluation efficiency and accuracy, we set the maximum output length to 8,192 tokens for all settings.
Appendix D Scaling to larger models
To validate the scalability of the MARSHAL framework and ensure our findings are not specific to the 4B parameter scale, we extended our training to the larger Qwen3-8B model. We trained an 8B-parameter MARSHAL generalist agent using the exact same set of strategic games (Tic-Tac-Toe, Kuhn Poker, and Mini Hanabi) and hyperparameters as the 4B experiments.
Strategic ability.
First, we evaluate the model’s proficiency on both training and held-out games. As shown in Table 5, the MARSHAL-8B agent achieves significant improvements over the Qwen3-8B base model across all environments. Notably, we observe the same strong generalization pattern seen in the 4B model: the agent generalizes effectively to held-out, complex games, achieving a drastic improvement in Leduc Hold’em (7.26% to 53.89%) and Simple Hanabi (4.55% to 37.27%).
Table 5: Strategic ability comparison on Qwen3-8B. The MARSHAL training yields consistent improvements on both training games and held-out testing games (Connect Four, Leduc Hold’em, Simple Hanabi).
Model Tic-Tac-Toe Kuhn Poker Mini Hanabi Connect Four Leduc Hold’em Simple Hanabi
Qwen3-8B 48.38 33.12 27.00 10.48 7.26 4.55
MARSHAL (Generalist, 8B) 54.05 44.49 55.28 21.55 53.89 37.27
Generalization to multi-agent systems.
We further evaluate whether these improved strategic capabilities translate to general reasoning benchmarks within multi-agent systems. We integrated the MARSHAL-8B agent into both the competitive MAD framework and the cooperative AutoGen framework.
The results, detailed in Table 6, confirm the scalability of our approach. The MARSHAL-trained 8B model consistently outperforms the base model:
  • In the competitive MAD setting, we observe an average improvement of 2.60%, with substantial gains on challenging math benchmarks like AIME (70.00% to 80.00%).
  • In the cooperative AutoGen setting, the improvement is even more pronounced, with an average gain of 3.90%, and a significant boost on AIME (60.00% to 70.00%).
These results demonstrate that the MARSHAL algorithm and game selection scale stably to larger models, consistently unlocking cooperative and competitive reasoning capabilities.
Table 6: Generalization to multi-agent systems using Qwen3-8B. The MARSHAL generalist consistently improves performance on reasoning benchmarks in both competitive (MAD) and cooperative (AutoGen) settings.
MAS Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MAD Qwen3-8B 82.49 95.00 96.36 83.46 70.00 90.00 89.59 53.03
MARSHAL (Generalist, 8B) 85.09 96.40 96.59 83.46 80.00 95.00 90.70 53.54
AutoGen Qwen3-8B 79.68 88.80 95.91 83.07 60.00 89.19 89.30 51.52
MARSHAL (Generalist, 8B) 83.58 94.40 95.00 85.04 70.00 95.00 90.04 55.56
Appendix E Comparison with SPIRAL: decoupling algorithm and game environments
To rigorously decouple the contributions of our proposed learning algorithm and our selection of game environments, we conducted a set of controlled experiments comparing MARSHAL against SPIRAL baseline methods. We analyze two specific variations:
  1. 1.
    Ablation on algorithm (RAE + our games): We train an agent using SPIRAL’s Role-conditioned Advantage Estimation (RAE) on our full selection of competitive and cooperative game environments. This isolates the impact of the advantage estimation technique.
  2. 2.
    Ablation on games (MARSHAL algorithm + competitive games): We train an agent using the MARSHAL algorithm but restricted to competitive-only games (Tic-Tac-Toe and Kuhn Poker), similar to SPIRAL. This isolates the impact of the cooperative games.
E.1 Strategic ability comparison
Table 7 presents the performance on both training and held-out games. We analyze the results through the lens of the two ablations:
Ablation on algorithm.
When replacing the MARSHAL algorithm with RAE (row 2), we observe a sharp performance decline in the cooperative games. Specifically, performance in Mini Hanabi falls from 54.33% to 24.93%, and in Simple Hanabi from 36.75% to 5.53%. This confirms that RAE struggles with the dense, multi-turn credit assignment required for cooperative settings, validating the necessity of MARSHAL’s turn-level "sum-then-normalize" technique to unlock the potential of self-play learning.
Ablation on games.
When training only on competitive games (row 3), the agent naturally fails to acquire cooperative skills (low scores in Hanabi). This serves as a crucial control baseline for analyzing the downstream transfer results below.
Table 7: Strategic ability comparison decoupling algorithm and game environments. MARSHAL (row 1) outperforms both the RAE-based agent (row 2) and the competitive-only agent (row 3), particularly in cooperative environments.
Model Tic-Tac-Toe Kuhn Poker Mini Hanabi Connect Four Leduc Hold’em Simple Hanabi
MARSHAL (Full Method) 53.13 40.05 54.33 19.98 54.56 36.75
SPIRAL (RAE) + Our Games 48.35 42.05 24.93 9.70 6.11 5.53
MARSHAL Alg. + Adv. Games 48.80 47.55 24.48 17.90 29.98 11.48
E.2 Generalization to multi-agent systems
We further evaluate how these variations affect transfer to general reasoning benchmarks within multi-agent systems.
Generalization to competitive systems (MAD).
As shown in Table 8, the full MARSHAL method achieves the highest average performance (75.96%). Notably, the RAE-based agent underperforms (74.23%), confirming the algorithmic superiority of MARSHAL. The competitive-only agent performs comparably to the full agent, which is expected as MAD is a competitive framework.
Generalization to cooperative systems (AutoGen).
Table 9 reveals the critical role of the game environments. The agent trained on the competitive-only games suffers a significant performance drop in the cooperative AutoGen framework (Average 80.34% vs. 82.15% for MARSHAL). This confirms that cooperative games are not merely "more data"; they are a necessary component for transferring skills to cooperative downstream tasks. Additionally, the RAE-based agent also underperforms, again solidifying the novelty of the algorithmic design of MARSHAL.
Table 8: Generalization to the competitive MAD framework.
Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MARSHAL (Full Method) 75.96 92.80 95.60 83.86 46.67 80.00 87.36 45.45
SPIRAL (RAE) + Our Games 74.23 92.40 96.06 84.25 36.67 77.50 86.79 45.96
MARSHAL Alg. + Adv. Games 75.24 92.20 95.60 83.46 46.67 80.00 86.88 41.92
Table 9: Generalization to the cooperative AutoGen framework. Note the drop in performance for the competitive-only games (row 3), highlighting the necessity of cooperative training.
Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MARSHAL (Full Method) 82.15 95.20 94.54 86.61 66.67 92.50 89.53 50.00
SPIRAL (RAE) + Our Games 80.39 94.80 94.77 86.61 60.00 92.50 88.57 45.45
MARSHAL Alg. + Adv. Games 80.34 94.20 94.24 86.61 60.00 90.00 89.37 47.98
Conclusion.
These decoupling experiments confirms the superiority of MARSHAL’s algorithmic design over SPIRAL’s RAE, and that MARSHAL’s successful generalization is not due to one factor alone. It requires both our selection of competitive and cooperative games to provide necessary learning signals, and our novel credit assignment algorithm to effectively learn from them.
Appendix F Comparison with MT-GRPO
A concurrent work MT-GRPO (zeng2025reinforcing) also identifies the limitations of trajectory-level estimation in the original GRPO and proposes a fine-grained turn-level strategy. While MT-GRPO shares our motivation, and is also a critical first step towards fine-grain credit assignment in multi-turn GRPO, our "sum-then-normalize" approach differs fundamentally from their strategy, offering distinct advantages in stability and scalability in the game tasks that we consider in this work.
F.1 Theoretical differences
MT-GRPO relies on normalizing the turn-level rewards turn-by-turn (i.e., computing the mean and variance for the reward at step $t$ across the batch). We identify two structural limitations with this approach that MARSHAL addresses:
  1. 1.
    The variable length problem: In real-world reasoning and gaming tasks, trajectory lengths vary significantly. With turn-by-turn normalization, the effective batch size shrinks as shorter trajectories conclude. This causes the variance of the normalization statistics to increase for later turns in the sequence, introducing instability for training. In contrast, MARSHAL normalizes the cumulative return, which remains defined for the entire trajectory regardless of length.
  2. 2.
    Immediate reward vs. long-term return: MT-GRPO’s local normalization focuses on the immediate step-wise reward relative to peers at that specific step. MARSHAL calculates the full Monte Carlo return first. By normalizing the cumulative outcome, we capture the long-term impact of actions relative to the global baseline, preserving critical long-term dependencies often obscured by local normalization.
F.2 Empirical Comparison
To validate this theoretical analysis, we implemented MT-GRPO and compared it directly against MARSHAL. We trained a Tic-Tac-Toe specialist using both algorithms and evaluated their strategic proficiency across our full game suite.
Strategic ability.
As shown in Table 10, the MT-GRPO agent exhibits notable performance drop across both the training game (Tic-Tac-Toe) and generalization to held-out games (e.g., Connect Four, Hanabi). In particular, the drop is significant in Mini Hanabi (50.48% to 36.67%), a game requiring consistent long-term coordination. This confirms that our "sum-then-normalize" formulation is more robust for the variable-length, multi-turn interactions inherent in strategic games.
Generalization to MAS.
We further extended this comparison to the integration within multi-agent systems to assess if the stability issues in game training affect general reasoning transfer. Table 11 present the results in the competitive MAD and cooperative AutoGen frameworks, respectively. MARSHAL consistently achieves higher average scores (75.01% vs. 73.77% in MAD; 80.15% vs. 79.10% in AutoGen) and outperforms MT-GRPO on key hard reasoning benchmarks like AIME and AMC. This indicates that the training proficiency provided by MARSHAL translates directly to better robust reasoning in general multi-agent tasks.
Table 10: Comparison between MARSHAL and MT-GRPO. The MARSHAL advantage estimation yields superior performance and generalization compared to the concurrent MT-GRPO method, validating the robustness of the "sum-then-normalize" approach.
Model (Tic-Tac-Toe Specialist) Tic-Tac-Toe Kuhn Poker Mini Hanabi Connect Four Leduc Hold’em Simple Hanabi
MARSHAL (Ours) 53.70 38.79 50.48 22.75 43.00 29.75
MT-GRPO 50.10 40.05 36.67 18.55 39.94 20.08
Table 11: Comparison of generalization to multi-agent systems between MARSHAL and MT-GRPO (using Tic-Tac-Toe specialists). MARSHAL achieves higher average performance and demonstrates greater stability in downstream tasks.
MAS Model (Tic-Tac-Toe Specialist) Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MAD MARSHAL (Ours) 75.01 92.20 96.06 83.07 43.33 82.50 86.76 41.12
MT-GRPO 73.77 92.60 95.91 84.65 36.67 77.50 86.91 42.13
AutoGen MARSHAL (Ours) 80.15 94.40 94.69 87.01 60.00 90.00 89.53 45.45
MT-GRPO 79.10 95.40 94.62 85.04 56.67 90.00 89.02 42.93
Appendix G Hyperparameter analysis on response length penalty
To ensure efficient and stable reasoning within the context window limits, our reward design includes a response length penalty characterized by three hyperparameters: the minimum length $l_{\min}$, the maximum threshold $l_{\max}$, and the penalty weight $\alpha$.
Selection of length constraints.
The length bounds were chosen based on practical engineering constraints:
  • $l_{\min}=11$: This is the hard lower bound determined by the minimum number of tokens required to generate a syntactically valid response structure with the correct format (e.g., <think></think><action>...</action>).
  • $l_{\max}=2048$: This soft upper bound is derived from the context window of the base model. Given Qwen3-4B’s 32k context length, we allocated a budget of approximately 16 turns per game history ($32,768/16=2048$), which provides a sufficient horizon for the reasoning traces in our chosen game portfolio while preventing context overflow.
Ablation on penalty weight $\alpha$.
To assess the sensitivity of the framework to the penalty weight $\alpha$, we conducted an ablation study training Tic-Tac-Toe specialists with $\alpha=1.0$ (stronger regularization), $\alpha=0.5$ (default), and $\alpha=0.0$ (no regularization). The results are detailed in Table 12.
Table 12: Ablation study on the length penalty weight $\alpha$. The default setting ($\alpha=0.5$) strikes a balance between performance and conciseness. Removing the penalty ($\alpha=0$) leads to overly verbose responses and performance degradation in cooperative games like Hanabi. "OL" denotes the percentage of games lost due to overlong responses.
Model Setup Strategic Performance Response Statistics
Tic-Tac-
Toe
Kuhn
Poker
Mini
Hanabi
Connect
Four
Leduc
Hold’em
Simple
Hanabi
Avg Len
OL in
Mini
Hanabi
OL in
Simple
Hanabi
$\alpha=0.5$ (Default) 53.70 38.79 50.48 22.75 43.00 29.75 1700 9.6% 14.3%
$\alpha=1.0$ (Strong) 54.90 42.82 47.53 18.25 50.46 26.05 1657 7.7% 13.9%
$\alpha=0.0$ (None) 54.65 40.95 38.18 23.35 28.61 20.10 1954 20.4% 27.3%
Analysis.
The results yield two key conclusions:
  1. 1.
    Robustness: The framework is robust to reasonable variations in $\alpha$. Increasing the regularization to $\alpha=1.0$ maintains comparable strategic performance to the default setting, indicating that the method is not overly sensitive to the precise magnitude of the penalty.
  2. 2.
    Necessity of regularization: The length penalty is crucial for stability. When removed ($\alpha=0$), the average response length increases significantly (from 1700 to 1954 tokens). This verbosity leads to a sharp performance drop in cooperative, theory-of-mind-demanding games like Mini Hanabi (50.48% to 38.18%). This degradation correlates directly with a dramatic increase in the "Overlong Response" failure rate (9.6% to 20.4%), confirming that the penalty effectively prevents the model from degenerating into computationally wasteful loops.
Appendix H Hyperparameters
To ensure reproducibility and consistency, we maintain a unified hyperparameter configuration for all trained models. The complete settings are provided in Table 13.
@@ Appendix H Hyperparameters
Appendix I Game observation and prompt
@@ 2. In each of your turns:
Hanabi
For the Hanabi variants, a cooperative imperfect information game, we provide the agent with a unique observation. An agent observes the cards held by all other players but remains unaware of its own hand. The observation also includes the number of remaining information and life tokens, the cards successfully played on the board (fireworks), and the contents of the discard pile. The prompt is highly structured, detailing the game rules, player-specific information, and strict response formatting. For training, we use Mini Hanabi, a simplified 2-player version with 2 colors, 2 ranks, 3-card hands, and 3 information/life tokens. For out-of-distribution evaluation, we use a more complex variant, Simple Hanabi, which features 3 colors, 2 ranks, 5-card hands, 8 information tokens and 3 life tokens. For example, an initial prompt details the starting setup of the chosen variant, including all visible cards and token counts, before asking the first player to make the first move (Listing 3).
@@
Appendix J Full results
In this section, we present the raw (unnormalized) results for all experiments in the main text.
Table 14: Full results of game-play return in training games
Model Tic-Tac-Toe Kuhn Poker Mini Hanabi
MCTS
(100 Simulations)
MCTS
(1000 Simulations)
@@
MARSHAL
   Tic-Tac-Toe 0.506/-0.358 -0.278/-0.379 -0.119/-0.142 2.019
   Kuhn Poker 0.397/-0.490 -0.388/-0.538 -0.107/-0.103 1.799
   Mini Hanabi 0.438/-0.644 -0.350/-0.702 -0.123/-0.130 2.222
   Generalist 0.544/-0.419 -0.302/-0.450 -0.114/-0.141 2.173
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe 0.760/-0.161 -0.378/-0.162 -0.146/-0.090 1.397
   Kuhn Poker -1.000/-1.000 -1.000/-1.000 -0.114/-0.117 0.000
   Mini Hanabi - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe 0.492/-0.517 -0.292/-0.567 -0.104/-0.091 1.392
   Kuhn Poker 0.267/-0.607 -0.426/-0.625 -0.074/-0.106 1.666
   Mini Hanabi 0.351/-0.788 -0.400/-0.811 -0.133/-0.084 2.128
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe 0.654/-0.376 -0.213/-0.366 -0.127/-0.126 1.764
   Kuhn Poker 0.391/-0.509 -0.446/-0.567 -0.116/-0.109 1.607
   Mini Hanabi 0.363/-0.732 -0.364/-0.764 -0.119/-0.128 2.100
@@
Model Connect Four Leduc Hold’em Simple Hanabi
MCTS
(10 Simulations)
MCTS
(100 Simulations)
@@
MARSHAL
   Tic-Tac-Toe -0.387/-0.703 -0.858/-0.974 -0.518/-0.702 1.785
   Kuhn Poker -0.448/-0.746 -0.874/-0.995 -0.460/-0.327 1.761
   Mini Hanabi -0.465/-0.885 -0.880/-0.998 -0.736/-0.403 2.036
   Generalist -0.421/-0.780 -0.871/-0.979 -0.483/-0.487 2.205
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe -0.593/-0.887 -0.939/-0.990 -0.632/-0.616 0.733
   Kuhn Poker -1.000/-1.000 -1.000/-1.000 -1.124/-1.003 0.000
   Mini Hanabi - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe -0.465/-0.754 -0.843/-0.989 -0.622/-0.553 1.143
   Kuhn Poker -0.408/-0.783 -0.899/-0.998 -0.789/-0.663 1.379
   Mini Hanabi -0.495/-0.939 -0.886/-0.998 -0.555/-0.483 1.841
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe -0.492/-0.790 -0.882/-0.983 -0.594/-0.460 1.303
   Kuhn Poker -0.460/-0.790 -0.908/-0.992 -0.753/-0.769 1.345
   Mini Hanabi -0.358/-0.898 -0.902/-0.998 -0.664/-0.389 1.925
@@
MARSHAL
   Tic-Tac-Toe 89.10 95.20 46.50 40.00 77.50 57.60 38.89
   Kuhn Poker 87.80 94.50 48.40 33.30 72.50 59.30 33.84
   Mini Hanabi 88.10 94.70 48.00 43.30 65.00 58.90 36.36
   Generalist 89.90 94.60 52.00 33.30 75.00 59.90 34.85
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe 89.10 94.80 53.90 50.00 72.50 63.00 37.88
   Kuhn Poker 86.20 94.00 45.70 33.30 72.50 53.80 33.84
   Mini Hanabi - - - - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe 88.80 94.80 48.80 40.00 72.50 57.70 36.36
   Kuhn Poker 88.60 94.00 48.40 40.00 72.50 57.90 34.34
   Mini Hanabi 87.80 94.50 48.40 36.70 70.00 59.30 36.87
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe 88.20 94.50 53.10 36.70 77.50 58.20 39.39
   Kuhn Poker 88.10 94.80 53.90 40.00 75.00 58.70 35.86
   Mini Hanabi 88.60 93.90 46.10 40.00 75.00 57.00 38.89
@@
MARSHAL
   Tic-Tac-Toe 92.20 96.06 83.07 43.33 82.50 86.76 41.12
   Kuhn Poker 91.60 96.21 82.68 40.00 82.50 87.39 41.41
   Mini Hanabi 91.40 95.60 82.68 43.33 77.50 87.04 38.38
   Generalist 92.80 95.60 83.86 46.67 80.00 87.36 45.45
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe 94.80 94.54 85.04 53.33 90.00 88.57 42.93
   Kuhn Poker 95.00 94.24 86.22 53.33 85.00 88.54 45.45
   Mini Hanabi - - - - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe 93.80 94.77 85.04 60.00 85.00 88.98 47.98
   Kuhn Poker 95.00 94.54 86.22 56.67 90.00 89.05 50.00
   Mini Hanabi 94.20 94.62 85.43 53.33 90.00 88.51 46.97
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe 94.40 94.62 84.65 56.67 87.50 89.33 46.97
   Kuhn Poker 95.00 94.69 84.25 56.67 85.00 89.62 47.47
   Mini Hanabi 94.20 94.47 87.01 63.33 90.00 88.92 43.94